### v1

@@ Abstract
keywords:
multiple dynamic networks, cluster analysis, spectral embedding
{MSCcodes}
60B20, 62H30
@@ 1 Introduction
A network is a powerful way to represent the relationships or connections among a collection of objects, and a dynamic network can further describe how the structure of these relationships or connections evolves over time, with important applications in many fields, including social network analysis [60, 53, 33, 41], neuroscience [19, 27, 21], and finance [18, 31, 51]. For example, [49, 13] study changepoint detection in a dynamic network whose evolution is governed by an underlying random walk process; [57, 2, 38] explore anomaly detection in single dynamic networks; [10, 9, 56] investigate community discovery in dynamic networks; and in [43, 50], the authors focus on clustering and online changepoint detection in diverse multilayer networks. In this paper, we also study dynamic networks, and explore the relationships among multiple dynamic networks, aiming to cluster them based on their evolutionary patterns.
@@
Brain neural networks provide a motivating example for the dynamic networks clustering problem: a practical problem in neuroscience is to identify synaptic connections that have significant influence during specific neural processes [32, 45]. For instance, during processes such as learning associations between stimuli and rewards or punishments, the evolution of brain neural networks can be studied through real experiments or simulations [19]. In [19], neuroscientists exploit techniques to simulate the time-varying behavior of brain neural networks by the removal of specific edges between specific pairs of neurons. For each edge, replicate simulations are conducted, resulting in multiple dynamic networks corresponding to the initial removal of a specific edge. We wish to analyze how dynamic networks evolve, and in particular whether the associated removed edge significantly impacts subsequent network development. If clustering reveals that the dynamic networks associated with a particular removed edge are highly similar to each other but significantly different from those of other edges, it suggests that this edge is important and has a potentially unique function in this process. Discerning patterns in network evolution is not restricted to neural networks. Related important inference questions arise in various domains, including the evolution of distinct subcommunities in social networks [41], interaction patterns in organizational networks [60], temporal changes in transportation networks [3], and anomalies in trade networks [50, 16, 12]. In all of these scenarios, a principled methodology for clustering networks based on evolutionary patterns would help address key application questions of interest.
@@
Due to the high dimensionality and sparsity common in network data, traditional clustering methods like k-means [4] are typically ineffective for accurately clustering dynamic networks by their distinct patterns of evolution. To address this challenge, we develop a dynamic network clustering method based on the Euclidean mirrors proposed in [8]. These mirrors constructed for individual dynamic networks are constructed using spectral decompositions of observed networks and provide a low-dimensional representation of network evolution. Spectral embedding methods are widely used and have proven effective for multi-sample network inference; see [34, 30, 6, 22, 29, 42, 3] for recent advancements. In [8], the authors assume that each node in the network has an associated time-varying low-dimensional latent vector of feature data, referred to as a latent position process, with the probabilities of the connections between any two nodes determined by their corresponding vectors, and the time-varying evolution of the latent vectors exhibiting a low-dimensional manifold structure. For a dynamic network, the differences between the latent position matrices at observed time points are measured pairwise, and then classical multidimensional scaling (CMDS) [11, 54, 35] is applied to obtain a configuration in a low-dimensional space that approximately preserves the dissimilarity. This results in a curve in low-dimensional Euclidean space that mirrors the evolution of the network dynamics over time, and is called the mirror of the dynamic network. Given the observed adjacency matrices of the network time series, the latent position matrices can be estimated using spectral embeddings [48, 47], with CMDS then providing an estimate of the mirror. The mirror method captures important features of the network evolution and has been demonstrated to be effective for visualizing a dynamic network and conducting inference tasks related to dynamic networks, including changepoint detection [8, 14, 13].
@@
For the problem of clustering evolutionary patterns of multiple dynamic networks, we consider the case where the differences in the evolution of the clusters can be reflected by the mirrors, and propose the Dynamic Network Clustering through Mirror Distance (DNCMD). The configuration obtained by CMDS is centered at the origin but has an inherent issue of identifiability concerning orthogonal transformations. Therefore, we construct a measure of the differences between estimated mirrors by solving an orthogonal Procrustes problem [44]. After obtaining the distances between the estimated mirrors of all pairs of dynamic networks, hierarchical clustering [39, 52] is applied to produce a dendrogram and the clustering of these associated mirrors based on the computed distances. We consider sufficiently distinct clusters evidence of distinct evolutionary patterns.
@@
For theoretical analysis, for a dynamic network, the evolution of the latent position matrix can be considered deterministic, as is the case when the features of the nodes in a network follow some predictable time-dependent pattern; it can also be random, as is the case when the actors in a network have underlying preferences that are subject to random shocks. [8] considers the latent position matrix to be random and assumes that all its rows, representing the latent positions of individual nodes, are generated from a stochastic process, and proves that the mirror of the dynamic network can be consistently estimated. In this paper, we prove the consistency result for the estimate of the mirror in the case where the latent position matrix is deterministic, allowing for flexibility in its evolution. Furthermore, for both cases, deterministic and random latent positions, we provide theoretical guarantees for DNCMD. We prove that DNCMD achieves exact recovery under mild conditions, where the term exact recovery means the clusters are correctly recovered with high probability.
@@
The structure of our paper is as follows. In Section 2, we introduce the model for multiple dynamic networks, where the clustering structure can be distinguished by the mirror, and in this section, we first consider the case where the latent position matrices are deterministic. In Section 3, we propose DNCMD to obtain the clustering result from the adjacency matrices at observed time points for multiple dynamic networks. In Section 4, we first provide the theoretical guarantee for DNCMD for deterministic latent positions, and then in Section 4.1, we briefly discuss the case where random latent positions are generated from stochastic processes, as considered in [8], and we build the model for multiple dynamic networks with a clustering structure and provide the theoretical guarantee for DNCMD in this scenario as well. Numerical simulations to demonstrate our theoretical results and show the clustering performance of DNCMD are presented in Section 5. Section 6 provides real data experiments for trade dynamic networks and the analysis of synaptic functions based on the evolution of brain neural networks. Detailed proofs of all stated results are presented in the supplementary material.
@@ 1.1 Notations
We summarize some notations used in this paper. For any positive integer $n$, we denote by $[n]$ the set $\{1,2,\dots,n\}$. For two non-negative sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we write $a_{n}\lesssim b_{n}$ ($a_{n}\gtrsim b_{n}$, resp.) if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$ ($a_{n}\geq Cb_{n}$, resp.) for all $n\geq 1$, and we write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. The notation $a_{n}\ll b_{n}$ ($a_{n}\gg b_{n}$, resp.) means that there exists some sufficiently small (large, resp.) constant $C>0$ such that $a_{n}\leq Cb_{n}$ ($a_{n}\geq Cb_{n}$, resp.). If $a_{n}/b_{n}$ stays bounded away from +$\infty$+ ∞, we write $a_{n}=O(b_{n})$ and $b_{n}=\Omega(a_{n})$, and we use the notation $a_{n}=\Theta(b_{n})$ to indicate that $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$. If $a_{n}/b_{n}\to 0$, we write $a_{n}=o(b_{n})$ and $b_{n}=\omega(a_{n})$. We say a sequence of events $\mathcal{A}_{n}$ holds with high probability if for any $c>0$ there exists a finite constant $n_{0}$ depending only on $c$ such that $\mathbb{P}(\mathcal{A}_{n})\geq 1-n^{-c}$ for all $n\geq n_{0}$. We write $a_{n}=O_{p}(b_{n})$ (resp. $a_{n}=o_{p}(b_{n})$) to denote that $a_{n}=O(b_{n})$ (resp. $a_{n}=o(b_{n})$) holds with high probability. We denote by $\mathcal{O}_{d}$ the set of $d\times d$ orthogonal matrices. Given a matrix $\mathbf{M}$, we denote its spectral, Frobenius, infinity norms, and the maximum entry (in modulus) by $\|\mathbf{M}\|$, $\|\mathbf{M}\|_{F}$, $\|\mathbf{M}\|_{\infty}$, and $\|\mathbf{M}\|_{\max}$, respectively.
@@ 2 Model
In this section, we introduce our model for multiple dynamic networks with a clustering structure. Suppose we have total $m$ dynamic networks $\{G_{t}^{(i)}\}_{i\in[m],t\in[T]}$ of $n$ common vertices for $T$ time points. We build the model for multiple dynamic networks on the framework for the single dynamic network detailed in [8]. Based on Random Dot Product Graph (RDPG), for each dynamic network $i\in[m]$, each vertex $s\in[n]$ at time $t\in[T]$ is associated with a latent position in $\mathbb{R}^{d}$, and edges between any two vertices arise independently with connection probability equal to the inner product of their respective latent position vectors. Therefore given latent position matrices $\{\mathbf{X}_{t}^{(i)}\in\mathbb{R}^{n\times d}\}_{i\in[m],t\in[T]}$, we have $G_{t}^{(i)}\sim\text{RDPG}(\mathbf{X}_{t}^{(i)})$ for all $i\in[m],t\in[T]$.
Definition 2.1 (Random dot product graph [7]).
@@
We call $\mathbf{P}=\mathbf{X}\mathbf{X}^{\top}$ the connection probability matrix. In this case, each $\mathbf{A}_{s,t}$ is marginally distributed as Bernoulli$(\mathbf{P}_{s,t})$ where $\mathbf{P}_{s,t}=\langle\mathbf{x}_{s},\mathbf{x}_{t}\rangle$.
RDPGs are a special case of latent position graphs or graphons [25, 17, 37]. In the general latent position graph model, each vertex $s$ is associated with a latent or unobserved vector $\mathbf{x}_{s}\in\mathcal{X}$ where $\mathcal{X}$ is some latent space such as $\mathbb{R}^{d}$, and given the collection of latent vectors $\{\mathbf{x}_{s}\}$, the edges are conditionally independent Bernoulli random variables with probability $\mathbf{P}_{s,t}=\kappa(\mathbf{x}_{s},\mathbf{x}_{t})$ for some kernel function $\kappa:\mathcal{X}\times\mathcal{X}\to[0,1]$. For RDPGs, $\kappa$ is the inner product.
Remark 2.2 (Orthogonal nonidentifiability in RDPGs).
@@
We first consider latent position matrices $\{\mathbf{X}^{(i)}_{t}\}$ as fixed to construct the model for simplicity and flexibility. We are also interested in the setting with randomly generated $\{\mathbf{X}^{(i)}_{t}\}$ as described in [8]. For example, the latent position of each vertex can be generated from a random process [13]. The model and theoretical results extended for random $\{\mathbf{X}^{(i)}_{t}\}$ are discussed in Section 4.1.
The key idea of [8] is to find a Euclidean analogue, called a mirror, which is a finite-dimensional curve that retains important signal of the evolution patterns in dynamic networks. More specifically, a mirror of a dynamic network of RDPG model with latent position matrix $\{\mathbf{X}_{t}\}_{t}$ is a dynamic vector $\{\mathbf{m}_{t}\}_{t}$ in a low-dimensional space, and the mirror keeps the overall trend of the dynamic network with that for any time points $t_{1}$ and $t_{2}$, the distance between $\mathbf{m}_{t_{1}}$ and $\mathbf{m}_{t_{2}}$ in the low-dimensional space is close to the “distance” between $\mathbf{X}_{t_{1}}$ and $\mathbf{X}_{t_{2}}$.
@@
For each fixed $i\in[m]$, consider the dynamic network $\{G^{(i)}_{t}\}_{t\in[T]}$ of $n$ vertices for $T$ time points and $G^{(i)}_{t}\sim\text{RDPG}(\mathbf{X}^{(i)}_{t})$ for all $t\in[T]$. Considering the orthogonal nonidentifiability of RDPGs as mentioned in Remark 2.2, for any pair of time points $t_{1}$ and $t_{2}$ in $[T]$, we define a distance to measure the dissimilarity between the two latent position matrices $\mathbf{X}^{(i)}_{t_{1}}$ and $\mathbf{X}^{(i)}_{t_{2}}$ as
(1) $\mathbf{D}^{(i)}_{t_{1},t_{2}}:=\frac{1}{\sqrt{n}}\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F},$
@@
for all $t_{1},t_{2}\in[T]$. Such low-rank configuration can be derived by classical multidimensional scaling (CMDS) [11, 54, 35] using the rank $r$ eigendecomposition of a matrix $\mathbf{B}^{(i)}$, which is obtained from $\mathbf{D}^{(i)}$ by double centering.
Definition 2.3 (CMDS embedding to dimension $r$).
@@
Remark 2.4 (Orthogonal nonidentifiability in CMDS embeddings).
@@
For the multiple dynamic networks clustering problem of interest, we suppose the total $m$ dynamic networks belong to $K$ cluster, and denote their cluster labels by $\{Y_{i}\}_{i\in[m]}$ where each label $Y_{i}\in[K]$. For the dynamic networks in the same cluster, we suppose they share the same distance matrix, i.e. for any $i,j\in[m]$ such that $Y_{i}=Y_{j}$ we have $\mathbf{D}^{(i)}=\mathbf{D}^{(j)}$, which means that they have similar change patterns of latent positions of the vertices across the time points. A special case is that the two dynamic networks share all the probability matrices across the time points and Theorem 2.5 shows that in this case the distance matrices for the latent positions are always identical. Notice that our model is flexible to describe situations as long as the latent positions of the vertices exhibit similar patterns of variation across the time points, and the case in Theorem 2.5 is just a special case.
Theorem 2.5.
@@
As we mentioned in Remark 2.4, the mirror is unique only up to an orthogonal transformation. We have Theorem 2.6 to describe the relationship between the mirrors of dynamic networks in the same clusters with the same distance matrix.
Theorem 2.6.
@@
Theorem 2.6 shows that in the same cluster, the dynamic networks have the same mirror up to orthogonal matrices. On the other hand, since CMDS, as a dimensionality reduction method, inevitably compresses some signals, it is theoretically possible to have the same mirror from different distance matrices. We are more interested in the situation where mirrors can differentiate between the evolution patterns corresponding to different clusters. We suppose there exists a distinct parameter $c>0$ such that for any two dynamic networks $i,j\in[m]$ with $Y_{i}\neq Y_{j}$, we have
(2) $\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}\geq c.$

### v2

@@ Abstract
footnotetext: This work was supported in part by Office of Naval Research (ONR) Science of Autonomy Award Number N00014-24-1-2278.
Keywords: multiple dynamic networks, cluster analysis, spectral embedding
@@ 1 Introduction
A network is a powerful way to represent the relationships or connections among a collection of objects, and a dynamic network can further describe how the structure of these relationships or connections evolves over time, with important applications in many fields, including social network analysis Zuzul et al. (2024); Wasserman and Faust (1994); Kumar et al. (2006); Palla et al. (2007), neuroscience Eschbach et al. (2020); Hutchison et al. (2013); Fox et al. (2005), and finance Ellington and Baruník (2020); Karim et al. (2022); Wanke et al. (2019). For example, Wang et al. (2021); Chen et al. (2024) study changepoint detection in a dynamic network whose evolution is governed by an underlying random walk process; Yu et al. (2018); Aggarwal et al. (2011); Manzoor et al. (2016) explore anomaly detection in single dynamic networks; Aynaud and Guillaume (2010); Aynaud et al. (2013); Xu and Hero (2014) investigate community discovery in dynamic networks; and in Pensky and Wang (2024); Wang et al. (2023), the authors focus on clustering and online changepoint detection in diverse multilayer networks. In this paper, we also study dynamic networks, and explore the relationships among multiple dynamic networks, aiming to cluster them based on their evolutionary patterns.
@@
Brain neural networks provide a motivating example for the dynamic networks clustering problem: a practical problem in neuroscience is to identify synaptic connections that have significant influence during specific neural processes Ko et al. (2011); Song et al. (2005). For instance, during processes such as learning associations between stimuli and rewards or punishments, the evolution of brain neural networks can be studied through real experiments or simulations Eschbach et al. (2020). In Eschbach et al. (2020), neuroscientists exploit techniques to simulate the time-varying behavior of brain neural networks by the removal of specific edges between specific pairs of neurons. For each edge, replicate simulations are conducted, resulting in multiple dynamic networks corresponding to the initial removal of a specific edge. We wish to analyze how dynamic networks evolve, and in particular whether the associated removed edge significantly impacts subsequent network development. If clustering reveals that the dynamic networks associated with a particular removed edge are highly similar to each other but significantly different from those of other edges, it suggests that this edge is important and has a potentially unique function in this process. Discerning patterns in network evolution is not restricted to neural networks. Related important inference questions arise in various domains, including the evolution of distinct subcommunities in social networks Palla et al. (2007), interaction patterns in organizational networks Zuzul et al. (2024), temporal changes in transportation networks Agterberg et al. (2022), and anomalies in trade networks Wang et al. (2023); De Benedictis et al. (2014); Chaney (2014). In all of these scenarios, a principled methodology for clustering networks based on evolutionary patterns would help address key application questions of interest.
@@
Due to the high dimensionality and sparsity common in network data, traditional clustering methods like k-means Ahmed et al. (2020) are typically ineffective for accurately clustering dynamic networks by their distinct patterns of evolution. To address this challenge, we develop a dynamic network clustering method based on the Euclidean mirrors proposed in Athreya et al. (2024). These mirrors constructed for individual dynamic networks are constructed using spectral decompositions of observed networks and provide a low-dimensional representation of network evolution. Spectral embedding methods are widely used and have proven effective for multi-sample network inference; see Levin et al. (2017); Jones and Rubin-Delanchy (2020); Arroyo et al. (2021); Gallagher et al. (2021); Jing et al. (2021); Pantazis et al. (2022); Agterberg et al. (2022) for recent advancements. In Athreya et al. (2024), the authors assume that each node in the network has an associated time-varying low-dimensional latent vector of feature data, referred to as a latent position process, with the probabilities of the connections between any two nodes determined by their corresponding vectors, and the time-varying evolution of the latent vectors exhibiting a low-dimensional manifold structure. For a dynamic network, the differences between the latent position matrices at observed time points are measured pairwise, and then classical multidimensional scaling (CMDS) Borg and Groenen (2007); Wickelmaier (2003); Li et al. (2020) is applied to obtain a configuration in a low-dimensional space that approximately preserves the dissimilarity. This results in a curve in low-dimensional Euclidean space that mirrors the evolution of the network dynamics over time, and is called the mirror of the dynamic network. Given the observed adjacency matrices of the network time series, the latent position matrices can be estimated using spectral embeddings Tang and Priebe (2018); Sussman et al. (2012), with CMDS then providing an estimate of the mirror. The mirror method captures important features of the network evolution and has been demonstrated to be effective for visualizing a dynamic network and conducting inference tasks related to dynamic networks, including changepoint detection Athreya et al. (2024); Chen et al. (2023, 2024).
@@
For the problem of clustering evolutionary patterns of multiple dynamic networks, we consider the case where the differences in the evolution of the clusters can be reflected by the mirrors, and propose the Dynamic Network Clustering through Mirror Distance (DNCMD). The configuration obtained by CMDS is centered at the origin but has an inherent issue of identifiability concerning orthogonal transformations. Therefore, we construct a measure of the differences between estimated mirrors by solving an orthogonal Procrustes problem Schönemann (1966). After obtaining the distances between the estimated mirrors of all pairs of dynamic networks, hierarchical clustering Nielsen and Nielsen (2016); Ward Jr (1963) is applied to produce a dendrogram and the clustering of these associated mirrors based on the computed distances.
DNCMD has been validated as effective for real-world complex problems, such as the neuroscience motivating example above. As shown in Figure 1, we analyze the dynamic evolution of neural networks after removing three different synaptic edge connections (see details in Section 6.1), and DNCMD successfully clusters most neural networks replicates for different removed connections and their impact on neural activity. This demonstrates its potential for identifying key synaptic connections, which we illustrate in Section 6.1.
Figure 1: Dendrogram of dynamic networks of neurons for the removal of each single edge, obtained using the DNCMD algorithm. For each removed edge we have $11$ replicates.
@@
For theoretical analysis, for a dynamic network, the evolution of the latent position matrix can be considered deterministic, as is the case when the features of the nodes in a network follow some predictable time-dependent pattern; it can also be random, as is the case when the actors in a network have underlying preferences that are subject to random shocks. Athreya et al. (2024) considers the latent position matrix to be random and assumes that all its rows, representing the latent positions of individual nodes, are generated from a stochastic process, and proves that the mirror of the dynamic network can be consistently estimated. In this paper, we prove the consistency result for the estimate of the mirror in the case where the latent position matrix is deterministic, allowing for flexibility in its evolution. Furthermore, for both cases, deterministic and random latent positions, we provide theoretical guarantees for DNCMD. We prove that DNCMD achieves exact recovery under mild conditions, where the term exact recovery means the clusters are correctly recovered with high probability.
@@
The structure of our paper is as follows. In Section 2, we introduce the model for multiple dynamic networks, where the clustering structure can be distinguished by the mirror, and in this section, we first consider the case where the latent position matrices are deterministic. In Section 3, we propose DNCMD to obtain the clustering result from the adjacency matrices at observed time points for multiple dynamic networks. In Section 4, we first provide the theoretical guarantee for DNCMD for deterministic latent positions, and then in Section 4.1, we briefly discuss the case where random latent positions are generated from stochastic processes, as considered in Athreya et al. (2024), and we build the model for multiple dynamic networks with a clustering structure and provide the theoretical guarantee for DNCMD in this scenario as well. Numerical simulations to demonstrate our theoretical results and show the clustering performance of DNCMD are presented in Section 5. Section 6 presents real data experiments to analyze synaptic functions in brain neural networks and temporal patterns in trade dynamic networks. Detailed proofs of stated results and additional experiment results are provided in the supplementary material.
@@ 1.1 Notations
We summarize some notations used in this paper. For any positive integer $n$, we denote by $[n]$ the set $\{1,2,\dots,n\}$. For two non-negative sequences $\{a_{n}\}_{n\geq 1}$ and $\{b_{n}\}_{n\geq 1}$, we write $a_{n}\lesssim b_{n}$ ($a_{n}\gtrsim b_{n}$, resp.) if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$ ($a_{n}\geq Cb_{n}$, resp.) for all $n\geq 1$, and we write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $a_{n}\gtrsim b_{n}$. The notation $a_{n}\ll b_{n}$ ($a_{n}\gg b_{n}$, resp.) means that there exists some sufficiently small (large, resp.) constant $C>0$ such that $a_{n}\leq Cb_{n}$ ($a_{n}\geq Cb_{n}$, resp.). If $a_{n}/b_{n}$ stays bounded away from $\infty$, we write $a_{n}=O(b_{n})$ and $b_{n}=\Omega(a_{n})$, and we use the notation $a_{n}=\Theta(b_{n})$ to indicate that $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$. If $a_{n}/b_{n}\to 0$, we write $a_{n}=o(b_{n})$ and $b_{n}=\omega(a_{n})$. We say a sequence of events $\mathcal{A}_{n}$ holds with high probability if for any $c>0$ there exists a finite constant $n_{0}$ depending only on $c$ such that $\mathbb{P}(\mathcal{A}_{n})\geq 1-n^{-c}$ for all $n\geq n_{0}$. We write $a_{n}=O_{p}(b_{n})$ (resp. $a_{n}=o_{p}(b_{n})$) to denote that $a_{n}=O(b_{n})$ (resp. $a_{n}=o(b_{n})$) holds with high probability. We denote by $\mathcal{O}_{d}$ the set of $d\times d$ orthogonal matrices. Given a matrix $\mathbf{M}$, we denote its spectral, Frobenius, infinity norms, and the maximum entry (in modulus) by $\|\mathbf{M}\|$, $\|\mathbf{M}\|_{F}$, $\|\mathbf{M}\|_{\infty}$, and $\|\mathbf{M}\|_{\max}$, respectively.
@@ 2 Model
In this section, we introduce our model for multiple dynamic networks with a clustering structure. Suppose we have total $m$ dynamic networks $\{G_{t}^{(i)}\}_{i\in[m],t\in[T]}$ of $n$ common vertices for $T$ time points. We build the model for multiple dynamic networks on the framework for the single dynamic network detailed in Athreya et al. (2024). Based on Random Dot Product Graph (RDPG), for each dynamic network $i\in[m]$, each vertex $s\in[n]$ at time $t\in[T]$ is associated with a latent position in $\mathbb{R}^{d}$, and edges between any two vertices arise independently with connection probability equal to the inner product of their respective latent position vectors. Therefore given latent position matrices $\{\mathbf{X}_{t}^{(i)}\in\mathbb{R}^{n\times d}\}_{i\in[m],t\in[T]}$, we have $G_{t}^{(i)}\sim\text{RDPG}(\mathbf{X}_{t}^{(i)})$ for all $i\in[m],t\in[T]$.
Definition 1 (Random dot product graph Athreya et al. (2018)).
@@
[…]
RDPGs are a special case of latent position graphs or graphons Hoff et al. (2002); Diaconis and Janson (2008); Lovász (2012). In the general latent position graph model, each vertex $s$ is associated with a latent or unobserved vector $\mathbf{x}_{s}\in\mathcal{X}$ where $\mathcal{X}$ is some latent space such as $\mathbb{R}^{d}$, and given the collection of latent vectors $\{\mathbf{x}_{s}\}$, the edges are conditionally independent Bernoulli random variables with probability $\mathbf{P}_{s,t}=\kappa(\mathbf{x}_{s},\mathbf{x}_{t})$ for some kernel function $\kappa:\mathcal{X}\times\mathcal{X}\to[0,1]$. For RDPGs, $\kappa$ is the inner product.
Remark 1 (Orthogonal nonidentifiability in RDPGs).
@@
We first consider latent position matrices $\{\mathbf{X}^{(i)}_{t}\}$ as fixed to construct the model for simplicity and flexibility. We are also interested in the setting with randomly generated $\{\mathbf{X}^{(i)}_{t}\}$ as described in Athreya et al. (2024). For example, the latent position of each vertex can be generated from a random process Chen et al. (2024). The model and theoretical results extended for random $\{\mathbf{X}^{(i)}_{t}\}$ are discussed in Section 4.1.
The key idea of Athreya et al. (2024) is to find a Euclidean analogue, called a mirror, which is a finite-dimensional curve that retains important signal of the evolution patterns in dynamic networks. More specifically, a mirror of a dynamic network of RDPG model with latent position matrix $\{\mathbf{X}_{t}\}_{t}$ is a dynamic vector $\{\mathbf{m}_{t}\}_{t}$ in a low-dimensional space, and the mirror keeps the overall trend of the dynamic network with that for any time points $t_{1}$ and $t_{2}$, the distance between $\mathbf{m}_{t_{1}}$ and $\mathbf{m}_{t_{2}}$ in the low-dimensional space is close to the “distance” between $\mathbf{X}_{t_{1}}$ and $\mathbf{X}_{t_{2}}$.
@@
For each fixed $i\in[m]$, consider the dynamic network $\{G^{(i)}_{t}\}_{t\in[T]}$ of $n$ vertices for $T$ time points and $G^{(i)}_{t}\sim\text{RDPG}(\mathbf{X}^{(i)}_{t})$ for all $t\in[T]$. Considering the orthogonal nonidentifiability of RDPGs as mentioned in Remark 1, for any pair of time points $t_{1}$ and $t_{2}$ in $[T]$, we define a distance to measure the dissimilarity between the two latent position matrices $\mathbf{X}^{(i)}_{t_{1}}$ and $\mathbf{X}^{(i)}_{t_{2}}$ as
$\mathbf{D}^{(i)}_{t_{1},t_{2}}:=\frac{1}{\sqrt{n}}\min_{\mathbf{O}\in\mathcal{O}_{d}}\|\mathbf{X}^{(i)}_{t_{1}}\mathbf{O}-\mathbf{X}^{(i)}_{t_{2}}\|_{F},$ (2.1)
@@
for all $t_{1},t_{2}\in[T]$. Such low-rank configuration can be derived by classical multidimensional scaling (CMDS) Borg and Groenen (2007); Wickelmaier (2003); Li et al. (2020) using the rank $r$ eigendecomposition of a matrix $\mathbf{B}^{(i)}$, which is obtained from $\mathbf{D}^{(i)}$ by double centering.
Definition 2 (CMDS embedding to dimension $r$).
@@
Remark 2 (Orthogonal nonidentifiability in CMDS embeddings).
@@
For the multiple dynamic networks clustering problem of interest, we suppose the total $m$ dynamic networks belong to $K$ cluster, and denote their cluster labels by $\{Y_{i}\}_{i\in[m]}$ where each label $Y_{i}\in[K]$. For the dynamic networks in the same cluster, we suppose they share the same distance matrix, i.e. for any $i,j\in[m]$ such that $Y_{i}=Y_{j}$ we have $\mathbf{D}^{(i)}=\mathbf{D}^{(j)}$, which means that they have similar change patterns of latent positions of the vertices across the time points. A special case is that the two dynamic networks share all the probability matrices across the time points and Theorem 1 shows that in this case the distance matrices for the latent positions are always identical. Notice that our model is flexible to describe situations as long as the latent positions of the vertices exhibit similar patterns of variation across the time points, and the case in Theorem 1 is just a special case.
Theorem 1.
@@
As we mentioned in Remark 2, the mirror is unique only up to an orthogonal transformation. We have Theorem 2 to describe the relationship between the mirrors of dynamic networks in the same clusters with the same distance matrix.
Theorem 2.
@@
Theorem 2 shows that in the same cluster, the dynamic networks have the same mirror up to orthogonal matrices. On the other hand, since CMDS, as a dimensionality reduction method, inevitably compresses some signals, it is theoretically possible to have the same mirror from different distance matrices. We are more interested in the situation where mirrors can differentiate between the evolution patterns corresponding to different clusters. We suppose there exists a distinct parameter $c>0$ such that for any two dynamic networks $i,j\in[m]$ with $Y_{i}\neq Y_{j}$, we have
$\frac{1}{\sqrt{T}}\min_{\mathbf{O}\in\mathcal{O}_{r}}\|\mathbf{M}^{(i)}\mathbf{O}-\mathbf{M}^{(j)}\|_{F}\geq c.$ (2.2)